# Internal acoustic resonances: small scattering center compared with the wavelength

### v1

@@ Abstract
In fields such as acoustics, electromagnetism, and quantum physics, the scattering of waves by localized objects is a fundamental phenomenon. Building on this, the present study investigates the energy distribution within a spherical scattering center during its interaction with an incident acoustic wave. The analysis reveals unexpected resonance effects driven by impedance mismatches between the scattering center and the surrounding ideal fluid. These resonance phenomena occur even when the magnitude order of the incident wavelength is significantly larger compared to the size of the scattering center. Such resonance behavior, observed in the small-particle regime, has potential applications in ultrasound imaging, targeted drug delivery, and metamaterial design.
@@ I Introduction
Wave scattering by localized objects (scattering centers) has been widely studied across fields such as acoustics, electromagnetism, classical mechanics, quantum physics, and particle theory [1]. The ongoing interest in scattering phenomena arises from their dual nature: the direct problem, which predicts scattering patterns based on the properties of the incident wave and the object, and the inverse problem, which determines the characteristics of the scattering centers from observed patterns [1]. This balance between prediction and deduction keeps scattering theory relevant in diverse applications. The origins of scattering research can be traced back to studies of acoustic phenomena, especially in naval engineering, where issues such as bursting air bubbles led to erosion and cavitation in propellers [2]. Rayleigh’s groundbreaking work on bubble dynamics [3] opened the door to significant advancements in the field, enabling later research into bubble vibration modes [4], acoustic luminescence [5], and broader applications in acoustic metamaterials [6, 7], medical diagnostics [8, 9], and virus inactivation [10]. These developments highlight the wide-reaching impact of acoustic scattering, from industrial to biomedical applications.
Building on these foundations, Anderson’s 1950 study on sound scattering by fluid spheres [11, 12] established the basis for research on more complex geometries, including cylinders [13], solid objects [14], and core-shell systems made from elastic and viscoelastic materials [15, 16, 17]. More recent findings reveal that materials like rubber silicones behave as fluid-like scattering centers due to their slow longitudinal sound propagation and negligible shear wave velocity when submerged in water [18, 19]. This behavior has practical implications for designing underwater isolation materials, with applications in submarines and naval vessels [20, 21]. In parallel with these practical advancements, theoretical progress has deepened our understanding of wave scattering across various media. The analysis of wave scattering involves key quantities such as energy and momentum transfer [22]. In elastic materials, energy storage is computed through the scalar product of displacement and traction vectors, while in fluids, pressure, and velocity waves are derived from scalar potentials, simplifying the analysis. Recent techniques proposed by Müller et al. [23] have drawn analogies between acoustic and electromagnetic scattering, expanding the theoretical framework and enhancing the analysis of wave interactions across multiple fields.
In contrast, acoustic scattering by small scattering centers (particles) compared with wavelength was often neglected due to their limited impact on direct wave propagation. Nevertheless, these particles can significantly affect energy transfer through mechanisms like Rayleigh scattering, which may either enhance or disrupt system performance [16, 24]. Accurately accounting for these effects is crucial, especially in applications requiring precise control, such as noise reduction, imaging technologies, and wave manipulation [25]. This study highlights the importance of addressing even minor scattering effects to improve overall system efficiency and reliability. Expanding the traditional concept of acoustic scattering, where an incident wave interacts with a scattering center, this research focuses on the dynamics of the wave that penetrates the scattering center. This study has already been addressed in optics with the finding of resonances and other related effects [26, 27, 28, 29, 30, 31].
This internal wave, influenced by the internal scattering coefficient, exhibits divergences under specific combinations of media and the size of the scattering center. Moreover, analytical expressions for energy transfer in the small-particle regime are derived and validated through practical examples using media such as air, petrol, and water. In one case, resonance is observed inside the scattering center due to a significant impedance mismatch between the medium and the scattering center. This article is structured as follows. In Sec. II, we present the scattering process of acoustic waves by a small fluid sphere in spherical geometry. The analytical expressions derived for the energy transferred during single scattering events. In Sec. III , we calculate the energy distribution within the fluid sphere for different media, such as air, petrol, and water. Resonances may occur even when the scatterer is much smaller than the wavelength of the incident wave, which is not expected. These resonance peaks, driven by the impedance mismatch between the scatterer and the surrounding fluid, lead to substantial energy storage, highlighting the importance of addressing even minor scattering effects to improve system efficiency and reliability Finally, in Sec. IV, our study provides a foundation for further investigations into more complex scattering geometries, such as core-shell structures, where interactions between materials could enhance energy localization and control.
Figure 1: Potential and kinetic energies as functions of the dimensionless size parameter $x_{0}$ for air (top) and petrol (bottom) in seawater. (top left) Small spherical scatterers resonate with incident acoustic waves at low impedance ratio, with potential energy dominating at low $x_{0}$ by several orders of magnitude. In contrast, (bottom left) the moderate impedance ration in petrol-seawater fail to support resonances entirely.
@@ II Acoustic wave scattering by a spherical particle
The propagation of harmonic acoustic waves in an ideal fluid with density $\rho_{0}$ is analyzed, following the formalism of Anderson and Faran [32, 14]. Since the acoustic waves are treated as small perturbations, the governing fluid equations are simplified through linearization. The wave propagates through the medium in an adiabatic manner, and the fluid is assumed to be irrotational. Under these conditions, both the pressure and velocity fields are derived from a scalar potential $\varphi(\mathbf{r})$ [33], which satisfies the Helmholtz equation $(\nabla^{2}+k^{2})\varphi(\mathbf{r})=0$, where $k$ is the wave number, related to the angular frequency $\omega$ by $k={\omega}/{c}$, with $c$ being the speed of sound in the fluid [34]. This scalar potential governs the spatial behavior of the perturbations in the homogeneous fluid. The pressure field, $p(\mathbf{r},t)=-\rho\partial_{t}\left[e^{-i\omega t}\varphi(\mathbf{r})\right]$, represents the compression and rarefaction of the fluid as the wave propagates. The velocity field, $\mathbf{v}(\mathbf{r},t)=e^{-i\omega t}\nabla\varphi(\mathbf{r})$, describes the oscillatory motion of fluid particles caused by the wave.
In addition, a spherical scattering center, which is also a fluid with radius $a$ and density $\rho_{1}$ (where $\rho_{1}\neq\rho_{0}$), is now incorporated into the system and is positioned at the origin of the coordinate system. The incident acoustic wave, represented by the potential $\varphi_{\text{inc}}(\mathbf{r})$, interacts with the scattering center as it propagates through the surrounding fluid. Inside the scattering center, the acoustic field is described by a scalar potential $\varphi_{1}(\mathbf{r})$, while the scattered wave outside is represented by $\varphi_{\text{sc}}(\mathbf{r})$. All the presented potentials satisfy the Helmholtz equation, as both the fluid and the scattering center are assumed to be acoustically isotropic, supporting only longitudinal compressional waves. At the surface of the scattering center, the boundary conditions ensure the continuity of physical quantities across the interface between the scattering center and the surrounding fluid, taking into account the chosen geometry [34]. These conditions are, at $r=a$,
@@
The first boundary condition ensures that the pressure inside the scattering center matches the pressure in the surrounding fluid at the surface $r=a$. The second boundary condition requires continuity of the normal component of the velocity field, ensuring that the velocity of fluid particles at the surface is consistent between the two fluids, inside and outside the scattering center[34, 33]. In this context, it is assumed that the incident wave is plane and monochromatic, so the incident wave equation is written as:
@@
In Eq. (2a), the incident wave $\varphi_{\text{inc}}$ is expanded in spherical harmonics, where $B$ represents the wave amplitude, and $k_{0}$ is the wave number in the surrounding fluid. The radial and angular components are captured by the spherical Bessel functions $j_{\ell}(z)$ and the Legendre polynomials $P_{\ell}(\cos\theta)$, respectively [35]. Since the wave inside the sphere retains the same modal decomposition as the incident wave, the internal wave in the scattering center is written similarly in Eq. (2b), with the coefficient $b_{\ell}$ representing the internal scattering strength. The spherical Bessel functions ensure regularity at the origin. Finally, Eq. (2c) describes the scattered wave, where spherical Hankel functions of the first kind, $h_{\ell}^{(1)}(k_{0}r)$, represent outgoing waves at large distances, and the scattering coefficient $s_{\ell}$ governs the amplitude of the scattered wave [35].
One important concept is acoustic impedance, $Z=\rho c$, that measures the resistance a medium offers to sound waves, where $c$ is the speed of sound.  The impedance ratio between the two media, $m_{t}=Z_{0}/Z_{1}$, plays a key role in determining how much of the incident wave is reflected or transmitted at the boundary. For simplicity, the parameters $x_{0}=k_{0}a$ and $x_{1}=k_{1}a=mx_{0}$ (where $m=c_{0}/c_{1}$) are defined. This simplification streamlines the notation, making it easier to calculate the coefficients $b_{\ell}$ and $s_{\ell}$ using the boundary conditions Eqs. (1a1b):
@@
Both coefficients share the same denominator, which may vanish for special combinations of $m$, $m_{t}$, and $x_{0}$. These divergences are also observed in far-field problems, particularly in the scattering field [36, 24].   Additionally, such a denominator can lead to resonance phenomena. At resonance, the system’s energy becomes significantly amplified due to constructive interference, making energy analysis a valuable tool for detecting and characterizing these phenomena [37]. Thus, the energy of a system composed of an acoustic wave interacting with a scattering center is
@@
The Eq. (4) characterizes the transfer of energy by sound waves through a medium. The potential energy arises from pressure variations induced by the sound wave, which alternately compresses and rarefies the fluid. The kinetic energy is associated with the oscillatory motion of fluid particles driven by the passage of the wave. Using the internal scalar potential $\varphi_{1}$ from Eq. (2b), it is possible to explicity calculate the potential and kinetic energy using Eq. (4). For the potential energy, the integral can be solved analytically yielding
$\frac{2f_{0}W_{K}}{Z_{1}\left|B\right|^{2}}$ $=x_{1}^{3}\sum_{\ell=0}^{\infty}\left|b_{\ell}\right|^{2}\left(2\ell+1\right)\left[\frac{\ell(2\ell+1)}{x_{1}^{3}}I_{1}(x_{1})-\frac{2\ell}{x_{1}^{3}}I_{2}(x_{1})+\frac{1}{2}\left(j_{\ell+1}^{2}\left(x_{1}\right)-j_{\ell}\left(x_{1}\right)j_{\ell+2}\left(x_{1}\right)\right)\right],$ (5a)
$\frac{2f_{0}W_{P}}{Z_{1}\left|B\right|^{2}}$ $=x_{1}^{3}\sum_{\ell=0}^{\infty}\left|b_{\ell}\right|^{2}\left(2\ell+1\right)\left[j_{\ell}^{2}\left(x_{1}\right)-j_{\ell-1}\left(x_{1}\right)j_{\ell+1}\left(x_{1}\right)\right].$ (5b)
with frequency $f_{0}$ of the incident wave, and $I_{1}\left(x_{1}\right)=\int_{0}^{x_{1}}j_{\ell}^{2}(x)\mathrm{d}x$, and $I_{2}\left(x_{1}\right)=\int_{0}^{x_{1}}j_{\ell}(x)j_{\ell+1}(x)\,x\,\mathrm{d}x$. The value $(2f_{0}/Z_{1}|B|^{2})$ normalizes the expressions for kinetic and potential energy, ensuring they are dimensionless. The integrals $I_{1}(x_{1})$ and $I_{2}(x_{1})$ lack an explicit analytical expression and thus must be calculated numerically. We truncate the summation of the partial-wave series at
@@
Acoustic impedance is a property that measures the opposition a medium offers to the transmission of sound waves. High acoustic impedance results in more resistance to sound wave propagation, while low acoustic impedance allows sound to pass more easily. A severe impedance mismatch disrupts this energy exchange, reducing the chances of resonance. Therefore, in systems with highly varied impedance, resonance effects are typically minimal or absent altogether.
@@ III.1 Systems with a impedance ratio $Z_{1}/Z_{0}\ll 1$
[…]
@@
Fig. 1 compares the potential and kinetic energies as functions of the dimensionless size parameter $x_{0}$ for an air bubble immersed in seawater. Fig. 1 (top left) shows the energy behavior for $\ell=0$, highlighting key resonance effects. Resonance occurs when an acoustic wave interacts with a scatterer at specific frequencies, causing constructive interference and a significant amplification of energy inside or around the scatterer. These resonance peaks represent points where the acoustic wave strongly interacts with the scatterer, leading to substantial energy deposition. Notably, a resonance peak appears at $x_{0}\ll 1$, corresponding to the monopole mode. This peak arises even though the scatterer is much smaller than the wavelength of the incident wave, primarily due to the impedance contrast between the air bubble and the surrounding seawater. Fig. 1 (top right) display the results for a sum over multiple values of $\ell$, truncated at the maximum value $\ell_{\text{max}}$ as described by Eq. (6). This truncation effectively captures the resonance behavior without increasing computational complexity. As $x_{0}$ increases, both curves exhibit oscillations related to resonance peaks. These oscillations become smoother and less pronounced over time, indicating that the system’s response is stabilizing.
@@
Similarly, the potential energy within the scatterer, as shown in Fig. 2, exhibits sharp resonance peaks as a function of $x_{0}$. For larger values of $x_{0}$, higher-order modes of $\ell$ begin to resonate, creating a comb-like structure of peaks. As $x_{0}$ increases further, the frequency of these peaks rises, indicating stronger resonance, as all modes interfere constructively when the scatterer approaches the wavelength size.
@@
Likewise, the kinetic energy follows a similar trend, as shown in Fig. 3.   Distinct resonance peaks appear as $x_{0}$ increases, with small-scale resonance clearly visible in the kinetic energy distribution, reflecting efficient energy transfer even for a small scatterer. As $x_{0}$ increases, the kinetic energy peaks across different values of $\ell$ become more synchronized. For $x_{0}>5$, the system reaches a regime where resonances align harmonically, suggesting strong and regular energy storage within the scatterer.
The existence of resonance behavior in Figs.1, 2 and 3 for certain values of $x_{0}$, representing combinations of the incident wavelength and particle radius, distinct peaks for both potential and kinetic energies are observed. These sharp peaks are unexpected in the small-particle regime, as resonance is typically associated with scattering centers of a size comparable to the incident wave.
$\frac{2f_{0}}{Z_{1}|B|^{2}}(W_{K})_{\text{Small}}$ $=\frac{m^{5}x_{0}^{5}}{45}\left[\frac{m_{t}}{m}\frac{3}{(3-m_{t}mx_{0}^{2})}\right]^{2},$ (7a)
$\frac{2f_{0}}{Z_{1}|B|^{2}}(W_{P})_{\text{Small}}$ $=\frac{4m^{3}x_{0}^{3}}{3}\left[\frac{m_{t}}{m}\frac{3}{(3-m_{t}mx_{0}^{2})}\right]^{2}.$ (7b)
The presence of a single resonance peak at $x_{0}\approx 0.014$ is revealed by Eq. (7), corresponding to the monopole mode ($\ell=0$). This peak demonstrates that significant energy can be stored within the scattering center, even when the scatterer is much smaller than the wavelength of the incident wave. This unexpected result is due to the strong acoustic impedance mismatch between the air bubble and the surrounding seawater.
The results highlight pronounced resonance behavior, particularly in the small-sphere regime. Both potential and kinetic energies exhibit sharp peaks at specific values of $x_{0}$, indicating significant resonance effects even when the size of the scattering center is much smaller than the wavelength of the incident wave. This challenges the usual expectation that energy storage would be negligible in such a regime. The analysis confirms that resonance effects can still lead to substantial energy being stored within the scattering center, due to the strong acoustic impedance mismatch between the air bubble and the surrounding medium. As $x_{0}$ increases, higher-mode resonances become more frequent, leading to increased energy storage across different modes. These findings may enable the development of acoustic metamaterials

### v2

# Trapped Acoustic Energy and Resonances in Spherical Scatterers

@@ Abstract
The effectiveness of biochemical antivirals are vulnerable to mutations, motivating physical approaches. Recent experiments with ultrasound reveal viral disruption at MHz frequencies, yet the mechanism remains unclear. We model viruses as fluid-like inclusions and analyze internal acoustic fields. Exact solutions reveal that impedance mismatch induces resonances even for subwavelength scatterers. Beyond the monopole, higher-order modes trap significant energy with significant resonance peak broadening even in absence of explicit dissipation mechanisms. These findings suggest internal acoustic resonances as a mechanism for viral destabilization and acoustic metamaterial applications.
@@ I Introduction
The fight against viral infections has long relied on biomolecular strategies [1], which typically employ specific binding proteins to trigger immune responses or block viral entry into host cells [2, 3]. Their effectiveness stems from the high specificity of binding to viral proteins [4], but this feature also makes them vulnerable to mutation-induced changes at the binding sites [5]. Physical approaches provide an alternative, since they target generic structural or mechanical properties that are less sensitive to mutations. Ultrasound, for instance, has shown potential to inactivate viruses by exciting resonant modes that compromise viral stability. Finite element simulations indicate that SARS-CoV-2 spike-protein bonds support torsional, bending, and rotational modes and can resonate in the 1–-20 MHz range [6]. In vitro experiments confirmed that SARS-CoV-2 envelopes rupture under ultrasound in the 5–-10 MHz range [7]. However, the experiments also demonstrated the rupture of other infectious viruses, including H1N1, reducing the chances of resonances tied to specific viral proteins.
Building on these observations and well-established models [8, 9, 10, 11, 12], we develop a theoretical framework for understanding ultrasound-driven destabilization of viral structures. Instead of modeling molecular bonds directly, we treat viruses as localized fluid inclusions and analyze the resulting acoustic scattering. This approach assumes that materials with negligible shear velocity behave as fluid-like scatterers [13, 14]. In addition, we also neglect dissipation in order to highligh the primary effects of acoustic resonances in biological media, following the seminal works of Epstein & Carhart and Allegra & Hawley, which have shown that losses primarily broaden the acoustic resonance without suppressing them [15, 16].
In acoustic scattering theory, small particles relative to the incident wavelength can strongly influence the energy transfer, with significant contributions arising from the monopole term. However, higher-order modes can build additional resonances on par with the ones from the monopole, storing higher internal energy storage and, thus, internal pressure or stress. These higher‑order modes are the acoustic counterparts of electromagnetic Mie resonances, which arise when an incoming field matches the natural modal frequencies of a sub‑wavelength scatterer [17, 18, 19, 20, 21, 22]. Such resonances concentrate energy and boost scattering cross‑sections far beyond Rayleigh’s long‑wavelength prediction [23], even with $ka\ll 1$ where the wavenumber $k=2\pi/\lambda$ and $a$ is the radius of the scattering center.
Here we argue that ultrasound-induced stresses, usually insufficient to disrupt 100–200 nm particles, are amplified by Mie resonances, enhancing internal fluid velocities and stress distributions, potentially destabilizing the structural integrity of viral scattering centers. By focusing in the internal energy stored inside the scattering centers, we derive closed‑form expressions for wave‑to‑scatterer energy transfer. Our findings reveal that the resonances are driven by the impedance mismatch between the scatterer and the surrounding fluid, suggesting the internal localization of acoustic waves characterized by a residence time $\tau_{\textrm{Wigner}}$, and the dynamical broadening of resonance peaks. We further demonstrate that higher-order modes amplify local energy density and scattering strength beyond a monopole-only picture. The article is structured as follows. In Sec.II, we present the scattering of acoustic waves by a small fluid sphere in spherical geometry and derive analytical expressions for the energy transferred in single scattering events. SectionIII reports numerical results for the internal energy distribution in various media for spherical scatterers. Finally, Sec.IV offers concluding remarks, relating Mie resonances to the phenomenon observed in Ref. [7].
@@ II Acoustic wave scattering by a spherical particle
Figure 1: Schematic of the scattering setup. A monochromatic acoustic plane wave ($\lambda_{0}$, $\mathbf{k_{0}}$) propagates along +z+z strikes a homogeneous sphere (radius $a$, density $\rho{1}$) embedded in a fluid of density $\rho_{0}$.
Figure 2: Potential and kinetic energies as functions of the dimensionless size parameter $x_{0}$ for air (left) and petrol (right) in seawater. Small spherical scatterers resonate with incident acoustic waves at low impedance ratio $\left(Z_{1}/Z_{0}\ll 1\right)$, with potential energy dominating at low $x_{0}$ by several orders of magnitude. In contrast, the moderate impedance ratio $\left(Z_{1}/Z_{0}\approx 1\right)$ in petrol-seawater fail to support resonances entirely.
The propagation of harmonic acoustic waves in an ideal irrotational fluid with density $\rho_{0}$ is analyzed [24, 11]. Under these conditions, perturbations in the fluid are described by the scalar potential $\varphi(\mathbf{r})$, which satisfy the Helmholtz equation $(\nabla^{2}+k_{0}^{2})\varphi(\mathbf{r})=0$, with wave number $k_{0}=\omega/c_{0}$, angular frequency $\omega$, and sound speed $c_{0}$  [25, 26]. The pressure field, $p(\mathbf{r},t)=-\rho\partial_{t}\left[e^{-i\omega t}\varphi(\mathbf{r})\right]$, represents the compression and rarefaction of the fluid as the wave propagates. The velocity field, $\mathbf{v}(\mathbf{r},t)=e^{-i\omega t}\nabla\varphi(\mathbf{r})$, describes the oscillatory motion of fluid particles caused by the wave.
In addition, a fluid-like spherical scattering center is positioned at the origin (see Fig. 1) with radius $a$, density $\rho_{1}$, sound speed $c_{1}$, and wavenumber $k_{1}=\omega/c_{1}$. The incident acoustic wave, represented by the potential $\varphi_{\text{inc}}(\mathbf{r})$, interacts with the scattering center as it propagates through the surrounding fluid. Inside the scattering center, the acoustic field is described by a scalar potential $\varphi_{1}(\mathbf{r})$, while the scattered wave outside is represented by $\varphi_{\text{sc}}(\mathbf{r})$. All the presented potentials satisfy the Helmholtz equation, with the following boundary conditions:
@@
For a monochromatic incident plane-wave, the solutions take the general form:
@@
Here B is the wave amplitude, $k_{0,1}$ is the wave number in the surrounding (internal) fluid; $P_{\ell}(\cos\theta)$ is the Legendre polynomial, $j_{\ell}(z)$ and $h_{\ell}(z)$ are the spherical Bessel and Henkel functions, respectively.
In terms of the dimensionless parameters $x_{0}=k_{0}a$ and $x_{1}=k_{1}a=mx_{0}$ ($m=c_{0}/c_{1}$), together with the boundary conditions, the absorption and scattering coefficients, $b_{\ell}$ and $s_{\ell}$, read:
@@
The impedance ratio $m_{t}=Z_{0}/Z_{1}=(\rho_{0}/\rho_{1})m$ governs wave transmission between media, with $m_{t}\gg 1$ leading to strong relfection ($b_{\ell}\rightarrow 0$) whereas $m_{t}\approx 1$ produces negligible scattering ($s_{\ell}\rightarrow 0$). In biological media, where $\rho_{1}\approx\rho_{0}$, transmission depends primarily on the medium organization and the impact on reducing the internal sound speed [27]. Both coefficients share the same denominator in Eq. (3), which may vanish for special combinations of $m$, $m_{t}$, and $x_{0}$. These divergences are also observed in far-field problems [28, 23].
The energy of a system composed of an acoustic wave interacting with a scattering center is
@@
The potential energy arises from pressure variations induced by the sound wave, while the kinetic energy is associated with the oscillatory motion of fluid particles driven by the passage of the wave. Using the internal scalar potential $\varphi_{1}$ from Eq. (2b), it is possible to explicitly calculate the potential and kinetic energy using Eq. (4) yielding
$\frac{2f_{0}W_{P}}{Z_{1}\left|B\right|^{2}}$ $=x_{1}^{3}\sum_{\ell=0}^{\infty}\left|b_{\ell}\right|^{2}\left(2\ell+1\right)\left[j_{\ell}^{2}\left(x_{1}\right)-j_{\ell-1}\left(x_{1}\right)j_{\ell+1}\left(x_{1}\right)\right].,$ (5a)
$\frac{2f_{0}W_{K}}{Z_{1}\left|B\right|^{2}}$ $=x_{1}^{3}\sum_{\ell=0}^{\infty}\left|b_{\ell}\right|^{2}\left(2\ell+1\right)\left[\frac{\ell(2\ell+1)}{x_{1}^{3}}I_{1}^{(\ell)}(x_{1})-\frac{2\ell}{x_{1}^{3}}I_{2}^{(\ell)}(x_{1})+\frac{1}{2}\left(j_{\ell+1}^{2}\left(x_{1}\right)-j_{\ell}\left(x_{1}\right)j_{\ell+2}\left(x_{1}\right)\right)\right].$ (5b)
with frequency $f_{0}$ of the incident wave. The value $(2f_{0}/Z_{1}|B|^{2})$ normalizes the expressions for kinetic and potential energy, ensuring they are dimensionless. The integrals $I_{1}^{(\ell}(x_{1})=\int_{0}^{x_{1}}j_{\ell}^{2}(x)\mathrm{d}x$ and $I_{2}^{(\ell}(x_{1})=\int_{0}^{x_{1}}j_{\ell}(x)j_{\ell+1}(x)\,x\,\mathrm{d}x$ lack an explicit analytical expression and thus must be calculated numerically. We truncate the summation of the partial-wave series at
@@ III.1 Systems with a impedance ratio $Z_{1}/Z_{0}\ll 1$
Acoustic impedance characterizes a medium’s resistance to the transmission of sound waves. A large impedance mismatch between two media results in a strongly reflective interface, thereby impeding the transfer of acoustic energy. This mechanism affects not only the incident wave but also waves generated within a finite scattering center and propagating toward the surrounding fluid. In the absence of dissipation, such waves can become effectively trapped inside the scatterer. Resonances arise when the rate of wave accumulation within the scattering center exceeds the rate in which energy is emitted into the surrounding medium [21].
The large difference in acoustic impedance between medium 0 and medium 1 significantly affects the energy distribution inside the scatterer. Here, we study the potential and kinetic energies within the scattering center, using Eqs. (5b) and (5a), the spherical scattering center was represented by an air bubble - medium 1 (density $\rho_{1}=1.205\,\text{kg/m}^{3}$ and speed of sound $c_{1}=343\,\text{m/s}$) - immersed in seawater - medium 0 (the density $\rho_{0}=1024\,\text{kg/m}^{3}$ and speed of sound $c_{0}=1522\,\text{m/s}$. The interaction between acoustic waves and the penetrable scatterer is examined as a function of the dimensionless size parameter $x_{0}$.
Figure 3: Potential energy distribution, as a function of the dimensionless size parameter $x_{0}$, for an air bubble immersed in seawater. Medium 0 corresponds to the surrounding seawater ($\rho_{0}=1024\text{kg/m}^{3}$, $c_{0}=1522\ \text{m/s}$), while medium 1 represents the air bubble ($\rho_{1}=1.205\text{kg/m}^{3}$, $c_{1}=343\,\text{m/s}$). Separate plots are shown for even (top) and odd (bottom) values of $\ell$. The resonance peaks, which grow in number as $x_{0}$ increases, indicate critical points where the acoustic wave is trapped and energy is deposited within the scattering center.
@@
Fig. 2 compares the potential and kinetic energies as functions of the dimensionless size parameter $x_{0}$ for an air bubble immersed (left) and a petrol bubble (right) in seawater. Fig. 2 (left) shows the energy behavior for $\ell=0$, highlighting key resonance effects. Resonance occurs when an acoustic wave interacts with a scatterer at specific frequencies, causing constructive interference and a significant amplification of energy inside or around the scatterer. These resonance peaks represent points where the acoustic wave strongly interacts with the scatterer, leading to substantial energy deposition. Notably, a resonance peak appears at $x_{0}\ll 1$, corresponding to the monopole mode. This peak arises even though the scatterer is much smaller than the wavelength of the incident wave, primarily due to the impedance contrast between the air bubble and the surrounding seawater. The resonances vanish for impedance ratio close to unity, as show in Fig. 2 (right).
The first monopole resonance, shown in Fig. 2, at $x_{0}=\sqrt{3/mm_{t}}\approx 0.014$, satisfies
$\frac{2f_{0}}{Z_{1}|B|^{2}}\left[W_{K}\right]_{\text{Small}}$ $=\frac{m^{5}x_{0}^{5}}{45}\left[\frac{m_{t}}{m}\frac{3}{(3-m_{t}mx_{0}^{2})}\right]^{2},$ (7a)
$\frac{2f_{0}}{Z_{1}|B|^{2}}\left[W_{P}\right]_{\text{Small}}$ $=\frac{4m^{3}x_{0}^{3}}{3}\left[\frac{m_{t}}{m}\frac{3}{(3-m_{t}mx_{0}^{2})}\right]^{2}.$ (7b)
@@
Similarly, the potential energy within the scatterer, as shown in Fig. 3, exhibits sharp resonance peaks as a function of $x_{0}$. For larger values of $x_{0}$, higher-order modes resonate, creating a comb-like structure. As $x_{0}$ increases further, the frequency of these peaks rises, as all modes interfere constructively when the scatterer approaches the wavelength size.
@@
Likewise, the kinetic energy follows a similar trend, as shown in Fig. 4. Distinct resonance peaks appear as $x_{0}$ increases, with small-scale resonance clearly visible in the kinetic energy distribution, reflecting efficient energy transfer even for a small scatterer. As $x_{0}$ increases, the kinetic energy peaks across different values of $\ell$ become far more synchronized and broader, with a characteristic linewidth $\Gamma$.
Resonances are classically described by $(E-E_{0}-i\Gamma/2)^{-1}$ with $\Gamma$ being associated as a damping mechanism. However, our model has no internal losses: this apparent damping is purely dynamical. Each cavity mode couples to the continuum of outgoing waves, so the surrounding fluid behaves like an absorbing wall at infinity. In Fano’s picture, a trapped mode interferes with a nonresonant background, turning ideal $\delta$ peaks into finite and often asymmetric profiles [20, 21, 30, 31]. Energy is not dissipated inside the inclusion; it is radiated after an average time $\tau_{\text{Wigner}}$.
In the diffraction limit ($ka\gg 1$), the solutions (3a) and (3b) simplify, yielding the $\mu$-th resonance position $\mathrm{X}_{\ell}^{\mu}=({\pi}/{2m})\left(2\mu+1+\ell\right)$. From this expression, the spacing between consecutive resonances is $\Delta_{\mu}X=\pi/m$ for fixed $\ell$, and $\Delta_{\ell}X=\pi/(2m)$ for fixed $\mu$. Since $\Gamma\sim 1/\tau_{\text{Wigner}}$ [20], it follows that $\tau_{\textrm{Wigner}}/\tau_{\textrm{min}}\sim(m/\pi)$, indicating that the wave remains trapped longer when $m\gg 1$ and minimum cycling time $\tau_{\textrm{min}}=2\pi a/c_{1}$. This regime is relevant for certain viral particles [32], where the contrast parameter can reach $m_{\textrm{HIV}}\approx o(10^{2})$.
Figure 5 plots the stored-energy contribution for all multipoles. Both kinetic and potential energy show that higher-order modes ($\ell>0$) have significant contributions even around resonances. Thus, a monopole-only approximation is asymptotically exact when $x_{0}\ll 1$; once $x_{0}\gg 0$, omitting $\ell>0$ underestimates the stored energy.
Figure 5: Relative modal contribution. Total potentail (black, dashed) and kinetic (red, full) energy in the small (right) and large (left) particle limit, highlighting the resonances. (inset) Although the monopole remains as the most significant mode overall, the energy histogram reveals a significant contribution from other modes.
The results highlight pronounced resonance behavior. Both energies show sharp peaks for some values of $x_{0}$, including for $x_{0}\ll 1$. The strong impedance mismatch between the medium and the scattering center leads to substantial internal energy storage. As $x_{0}$ increases, resonances become denser and other modes contributions cannot be neglacted.